OpenStudy (anonymous):

Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x².

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@melchar

jimthompson5910 (jim_thompson5910):

I'll write one example to get you started \[\Large x^{1/2}*x^{3/2} = x^{1/2+3/2}\] \[\Large x^{1/2}*x^{3/2} = x^{(1+3)/2}\] \[\Large x^{1/2}*x^{3/2} = x^{4/2}\] \[\Large x^{1/2}*x^{3/2} = x^{2}\]

OpenStudy (anonymous):

which property is this

OpenStudy (anonymous):

@jim_thompson5910